You stop by the union to grab a chick-fil-a sandwich (440 Calories), waffle fries (400 Calories) and a 16 oz pepsi (200 Calories). If you burn 74.0 Calories per mile of walking, how far would you need to walk to burn off your meal?
Answer #1

Calories which i need to burn = 440+400+200 = 1040 calories

to burn 74 calories ,i need to walk 1 mile.

so to burn burn off my meal , i need to walk = (1/74) x 1040 = 14.05 mile . Answer